the set of big bang theory needs to be repainted .one gallon of paint covers450 square feet.the set is [24 feet] 24 feet by 18 feet and the wall are 12 feet high.how many full gallons of paint will be needed to paint the set.

Answer by checkley79(3341)

I'll assume one of the 24 feet figures is redundant.
24*18=432 ft^2 for the floor.
2(24*12)
2*288=576 ft^2 for the 2 larger walls.
2(18*12)
2*216=432 ft^2 for the 2 smaller walls.
432+576+432=1,440 ft^2 for the surface area.
1,440/450=3.2 gallons of paint is required.